A 140 kN force with a slope of 4 vertical and 7 horizontal is located on the second quadrant. Determine its y- component.

1
Expert's answer
2022-03-28T11:02:34-0400
"F_y=140\:\rm N\times \frac{4}{\sqrt{4^2+7^2}}=57\:\rm N"